OpenStudy (anonymous):

The 8th term of an arithmetic sequence is 11 and the 15th term is 21. Find the common difference, the first term of the series and the nth term. The 8th term of an arithmetic sequence is 11 and the 15th term is 21. Find the common difference, the first term of the series and the nth term. @Mathematics

OpenStudy (anonymous):

a1=1,d=10/7,an=1+(10/7)(n-1)

OpenStudy (anonymous):

can the common difference be a fraction ?

OpenStudy (anonymous):

coud be anything

OpenStudy (anonymous):

don't u wanna know how i got these answers...

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

i know how to do it, i wasnt sure if my common difference was right, because i got a fraction..

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

thanks a lot ! :)